OpenStudy (anonymous):

Combine the like terms to simplify the expression: 4 + 7xy2 + 5y3 + 10x –4xy2 – 8x – 4y3 + 9 A. y6 + 3x2y2 + 10x + 13 B. y3 + 3xy2 + 2x + 13 C. y3 + 3xy2 + 10x + 13 D. y3 – 3xy2 + 2x + 9

OpenStudy (anonymous):

B @200205650

OpenStudy (anonymous):

Gather the like terms first. \[4+(7x{y}^{2}-4x{y}^{2})+(5{y}^{3}-4{y}^{3})+(10x-8x)+9\] From there on, simplify.

OpenStudy (anonymous):

constants: 4+9 = 13 xy^2: 7-4 = 3xy^2 y^3: 5-4 = y^3 x: 10-8 = 2x y3 + 3xy2 +2x +13 (answer D)

OpenStudy (anonymous):

sorry B

OpenStudy (anonymous):

constants: 4+9 = 13 xy^2: 7-4 = 3xy^2 y^3: 5-4 = y^3 x: 10-8 = 2x y3 + 3xy2 +2x +13 (answer D)

OpenStudy (anonymous):

ok?????